# MaskCLIP++: A Mask-Based CLIP Fine-tuning Framework for Open-Vocabulary Image Segmentation

### v2

@@ Appendix C Additional Experiments
C.1 Additional Illustrations
Implementation details of Figure 1 We use mask generation and mask classification as the inference modes, with Figure 1(a) and Figure 1(b) presenting the performance upper bounds of the ideal mask generator and ideal mask classifier, respectively. The ideal mask generator is implemented by directly using the ground truth masks. To implement the ideal mask classifier, we compute the maximum matching between the generated masks and ground truth masks based on IoU, then assign the corresponding GT categories to the matched masks and combine them into the output segmentation maps.
Figure 5: A toy example: Because the inconsistency alignment between $s$ (similarity before $\Theta$) and $r$ (similarity after $\Theta$), after a round of gradient descent, $s$ is further from the target $\hat{s}$.
Phenomenon of inconsistency alignment. We first use a toy example to illustrate the potential consequences of destroying alignment consistency. Assume that we use the parameter $\Theta$ to update the mask embeddings, changing the original similarity from $s$ to $r$, and then optimize towards the target. However, due to the lack of explicit constraints on $\Theta$, there may be a situation as shown in Figure 5, where the order of $s$ and $r$ is inconsistent, resulting in correct optimization on $r$ but incorrect optimization on $s$. For simplicity, $\mathcal{L}$ in Figure 5 denotes an L1 loss without softmax, and only mask embeddings $m$ are updated in one round of gradient descent.
Futuremore, we use Table 12 as a supplement to Table 2 to illustrate the extent to which different PSM structures destroy the consistency alignment constraint. For models fine-tuned with various PSMs, we test by retaining or removing the PSM module, and compare the performance differences to assess whether the alignment consistency is maintained before and after the PSM intervention. The results show that only the three blue PSM structures largely satisfy the consistency alignment constraint, while the three red PSM structures severely disrupt it.
Table 11: The phenomenon of consistency alignment.
PSM test mIoU(S) mIoU(U) mIoU
- - 32.7 16.3 23.3
$\langle\operatorname{Linear}(E_{m}),E_{t}\rangle$ 37.7 20.2 27.7
32.4 1.9 14.9
$\langle E_{m},\operatorname{Linear}(E_{t})\rangle$ 37.6 20.3 27.6
32.5 1.8 14.9
$\operatorname{Linear}\langle E_{m},E_{t}\rangle$ 45.2 24.6 33.4
45.4 25.7 34.1
$\langle E_{m},\operatorname{Norm}(P_{t})\rangle$ 39.5 21.4 29.1
31.9 4.5 16.2
$\langle E_{m},\operatorname{Norm}(E_{t})\rangle$ 45.8 24.4 33.5
46.2 25.2 34.1
$\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle$ 44.0 24.1 32.6
46.3 26.7 35.1
Table 12: Performance of MaskCLIP++ on closed-vocabulary image segmentation tasks.
CLIP arch. Dataset mIoU PQ AP
ConvNeXt-B COCO-Stuff 47.1 - -
ViT-L/14 COCO-Stuff 51.0 - -
ConvNeXt-B COCO-Panoptic 57.6 46.1 34.4
ViT-L/14 COCO-Panoptic 63.3 52.3 42.3
The performance of closed-vocabulary image segmentation. We present the performance of MaskCLIP++ on closed-vocabulary image segmentation in Table 12, with the same configuration as in Table 5 and Table 6, respectively.
C.2 Additional Ablations
(a) ConvNeXt-B
(b) ConvNeXt-L
(c) ViT-B/16
(d) ViT-L/14
Figure 6: Under various input sizes of different CLIP models, our method (blue bars) has significant improvements over the original CLIP (red bars). The input size of the mask generator remains unchanged. All models are fine-tuned on COCO Panoptic and evaluated on ADE20K with mIoU reported.
Impact of CLIP input size. Considering that input image size often has a significant impact on the mIoU of semantic segmentation models, we compared the performance of our method with the original CLIP under different input sizes. As shown in Figure 6, with a fixed input size for the mask generator, our fine-tuned CLIP consistently outperforms the original CLIP across various input sizes for open-vocabulary semantic segmentation.
Overall, segmentation performance initially increases and then decreases as input size grows. Since our fine-tuned CLIP can better recognize local image features, the optimal input size tends to be smaller than that of the original CLIP. Our framework allows us to maintain the input size for the mask generator while reducing the input size for mask recognition, thereby enhancing performance and reducing computational costs.
Impact of integrating image context information into mask representation. Averaging CLIP features over perfectly precise mask regions is not conducive to extracting mask representations with stronger recognition capabilities, as recognition tasks typically require contextual information and focused attention on target internal regions. Here, we distinguish how we address this common issue using CLIP’s original pooling method. For CLIP models utilizing Attention Pooling (e.g., ViT-B/16), the hyperparameter $\alpha$ controls the degree of contextual information incorporated in the mask representation. As shown in Table 1313(b), smaller $\alpha$ values, indicating more extensive use of contextual information, lead to better performance. For CLIP models employing global average pooling (e.g. ConvNeXt-B), as shown in Table 1313(a), using global image embeddings $E_{i}$ or patch embeddings $E_{p}$ to construct pseudo-text embeddings $P_{t}$ in the PSM stage enhances the utilization of image context information to some extent, thereby improving performance.
Choice of Loss. Table 15 presents the performance differences observed when fine-tuning CLIP ConvNeXt-B with various loss functions. Training with COCO Panoptic (Lin et al., 2014) and evaluating on ADE20K (Zhou et al., 2017). When the training labels for a single image include multiple instances of the same category, such as pixels or instance masks, we re-weight the loss by categories to prevent bias toward more common classes during optimization. We adopt a class-reweighted cross-entropy loss for all our experiments in the main text.
Position of the mask prior insertion. Due to the repetitive nature of the ViT architecture, the position of mask prior insertion in CLIP models based on such architectures is not absolute. Following the approach of Xu et al. (2023c) for constructing attention masks, we can insert the mask prior at an earlier position than the last attention block. Table 15 takes ViT-B/16 as an example, demonstrating that inserting the mask prior at the third-to-last layer yields better performance. Therefore, in Table 4, the mask prior is inserted at the third-to-last layer for all ViT-based CLIP models.
Table 13: Impact of integrating image context information into mask embeddings $E_{m}$. The implementation varies based on the original CLIP’s pooling method.
(a) The case of CLIP using global average pooling, taking ConvNeXt-B as an example. $\Psi$ is a placeholder for image context information, and $P_{t}$ defined in equation 3 can be rewritten as $P_{t}=\operatorname{Attn}(q(E_{t}),k(\Psi),v(\Psi))$. $P_{t}$ interacts with $E_{m}$ in the PSM stage to increase the utilization of image context information.
Method in PSM Image Context mIoU
No $\Phi$ 34.1
$\Phi=E_{i}$ 34.5
$\Phi=E_{p}$ 35.1
(b) The case of CLIP using attention pooling, taking ViT-B/16 as an example. $\alpha$ is defined in equation 2. The bigger $\alpha$ is, the less the semantic similarity of the image context is utilized. The learnable $\alpha$ is initialized from $e^{-5}$ and learned in the logarithmic space.
Method in CLIP Image Context mIoU
$\alpha=e^{5}$ 31.1
$\alpha=e^{-5}$ 33.5
learnable $\alpha$ 33.8
Table 14: Choice of Loss. “Re.” means the loss is reweighted by categories. CLIP ConvNeXt-B is used.
Prior Loss Re. mIoU(S) mIoU(U) mIoU
Pixel CE 39.5 21.2 29.0
CE 44.3 23.4 32.8
Instance Mask BCE 45.7 24.5 33.5
BCE 46.3 25.3 34.3
CE 45.5 26.0 34.4
CE 46.3 26.7 35.1
Table 15: The impact of the position of mask prior insertion for ViT-B/16 CLIP. “Prior pos.” denotes the insert position relative to the total number of attention blocks.
Prior pos. mIoU PQ
-1 33.3 24.1
-2 33.5 24.5
-3 33.8 24.4
-4 33.5 24.4
Qualitative results. Figure 7 compares the visual results of open-vocabulary semantic segmentation using different methods with the same mask generator. We show more qualitative results of semantic- and instance level of open-vocabulary segmentation in Figure 8.
Image                    CLIP                MAFT+              MaskCLIP++ GT
Figure 7: Visualizations of open-vocabulary semantic segmentation on ADE20K (Zhou et al., 2017). From left to right are the input image, the result using the original CLIP, the result of MAFT+ (Jiao et al., 2024), the result using MaskCLIP++, and the ground truth. The same mask generator is used for CLIP, MAFT+ and MaskCLIP++.
@@ C.2 Additional Ablations
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 8: More visualizations of semantic- and instance-level of open-vocabulary segmentation. We use mask generator from FC-CLIP (Yu et al., 2024) with ConvNeXt-L backbone. And use CLIP ViT-L/14 fine-tuned on COCO Panoptic (Lin et al., 2014). Images are from ADE20K (Zhou et al., 2017). The confidience threshold for instance segmentation is 0.3.

### v3

@@ Appendix B Dataset Details
Method BDD100K Dark Zurich MHPv1 FoodSeg103 ATLANTIS DRAM
OpenAI CLIP 44.6 27.5 48.6 57.1 58.1 89.3
EVA02 CLIP 44.7 34.3 49.2 61.4 61.0 91.0
CLIPSelf 51.8 30.6 61.0 47.4 59.1 65.3
CAT-Seg 59.3 40.3 49.8 54.6 65.3 94.6
MaskCLIP++ 67.2 48.3 64.9 64.7 64.5 98.0
(a) General
Method iSAID ISPRS Potsdam WorldFloods FloodNet UAVid
OpenAI CLIP 59.2 65.5 65.9 65.9 48.9
EVA02 CLIP 61.0 68.4 71.4 59.6 42.2
CLIPSelf 43.6 41.7 42.4 38.3 62.6
CAT-Seg 59.8 76.9 71.0 62.3 64.4
MaskCLIP++ 60.5 80.0 75.1 63.0 64.4
(b) Earth Monitoring
SUIM CUB-200 CWFID
55.5 54.7 42.8
57.6 60.0 39.7
56.0 51.1 61.9
72.9 47.2 73.0
73.2 40.8 82.5
(c) Agriculture and Biology
Method Kvasir Inst. CHASE DB1 CryoNuSeg PAXRay-4
OpenAI CLIP 93.2 50.0 50.0 50.8
EVA02 CLIP 95.0 50.0 50.0 51.7
CLIPSelf 72.0 50.0 50.0 50.0
CAT-Seg 89.8 50.0 50.0 58.1
MaskCLIP++ 99.2 50.0 50.0 99.4
(d) Medical Sciences
Corrosion CS DeepCrack ZeroWaste-f PST900
25.5 69.2 32.5 19.4
25.0 58.9 33.0 18.4
31.6 76.8 45.2 27.5
21.6 61.0 30.2 26.7
24.4 85.7 57.4 33.8
(e) Engineering
Table 12: Mask accuracy on MESS datasets [4].
@@ Appendix C Additional Experiments
C.1 Mask accuracy on MESS
Tab. 12 complements Tab. 1 by listing the mask accuracy of different models on each dataset in MESS [4]. MaskCLIP++ achieves better maskAcc on 17 out of 22 datasets, providing compelling empirical evidence that the proposed method significantly enhances mask classification accuracy while simultaneously preserving the inherent generalization capabilities of the original CLIP.
However, note that in specialized fine-grained classification tasks like the bird-oriented CUB-200 dataset [58], where category labels are semantically close and images usually have a single foreground subject, fine-grained category domain-specific knowledge impacts performance more than localized perception. Thus, CLIP models fine-tuned on COCO generally perform worse than the original CLIP in such cases.
C.2 Effectiveness of mask tuning on different CLIP architectures
To validate the generalizability of our approach across varying CLIP architectures, we conduct fine-tuning experiments on the COCO Panoptic dataset [36] using distinct CLIP architectures and report the mask classification accuracy of our method as documented in Tab. 13.
Arch. A-847 PC-459 A-150 PC-59 Stuff
RN50x16 24.1 37.5 50.4 73.3 54.1
ConvNeXt-B 30.0 42.1 56.4 76.3 55.6
ConvNeXt-L 33.0 50.8 61.0 78.7 59.0
ConvNeXt-XXL 38.7 56.1 64.6 81.5 62.5
ViT-B/16 29.3 45.7 58.3 78.6 58.0
ViT-L/14 37.5 56.9 66.0 83.2 63.6
Table 13: Mask accuracy of MaskCLIP++ on different architectures fine-tuned on COCO Panoptic [36].
C.3 Other ablation studies
Effect of Fuser depth. The positioning of mask insertions exerts non-negligible effects on model performance. As delineated in Sec. 3.2, the quantity of Transformer blocks in the Fuser module can be mathematically formulated as $L-K$. Tab. 14(a) presents a systematic analysis of how varying Fuser depths impact OVS performance. Considering the computational overhead that scales linearly with mask quantity when increasing Fuser depth, we empirically select $L-K=2$ as the default configuration.
$\bm{L-K}$ mIoU PQ
1 33.3 24.1
2 33.5 24.5
3 33.8 24.4
4 33.5 24.4
(a) Effect of Fuser depth.
$\bm{\alpha}$ mIoU
$e^{5}$ 33.1
$e^{-5}$ 33.5
learnable 33.8
(b) Effect of $\alpha$.
Table 14: Ablation studies on the effect of Fuser depth and $\alpha$.
Effect of $\alpha$. As defined in Eq. 2, the bigger $\alpha$ is, the less the semantic similarity of the image context is utilized. As $\alpha\rightarrow+\infty$, $E_{m}^{(l+1)}$ degenerates to the mask average pooling result of $F^{(l)}$, thereby inadequately leveraging contextual information beyond the masked regions. Conversely, when $\alpha=0$, $E_{m}^{(l+1)}$ reduces to the CLS token after $l$-th layer. As shown in Tab. 14(b), selecting a smaller $\alpha$ value yields superior performance. Therefore, we initialize $\alpha$ as $e^{-5}$ and adaptively adjust it in the logarithmic space during training.
Effect of classification ensemble. To demonstrate that the performance gains of our method primarily stem from enhanced mask classification capabilities, we present in Tab. 15 the comparative analysis of implementing versus omitting the ensemble strategy.
Method Ensemble A-847 PC-459 A-150 PC-59
FC-CLIP 12.1 12.7 27.4 42.8
14.8 18.2 34.1 58.4
MaskCLIP++ 15.3 21.4 36.8 62.4
15.4 21.3 37.1 62.6
Table 15: Effect of classification ensemble.
@@ C.3 Other ablation studies
Figure 7: More visualizations of semantic- and instance-level of open-vocabulary segmentation. We use mask generator from FC-CLIP [76] with ConvNeXt-L backbone. And use CLIP ViT-L/14 fine-tuned on COCO Panoptic [36]. Images are from ADE20K [80]. The confidience threshold for instance segmentation is 0.3.